OpenStudy (anonymous):

Can someone please explain how to do this? I have sevral questions like this: 3x - 2y = 12 find the x-intercept (a) 6, (b) 4, (c) -4, (d) -6

OpenStudy (hba):

For x-intercept plug in y=0

OpenStudy (anonymous):

So, the answer is 4?

OpenStudy (hba):

3x-2(0)=12 3x=12 x=12/3 x=4 :) Yeah right. Welcome to Openstudy :)

OpenStudy (anonymous):

thank you

OpenStudy (hba):

Welcome :)